Hello I need help with this problem. Thank you.
Problem:
a/24 = 15/16

Answer by stanbon(75887)

a/24 = 15/16
----
Multiply both sides by 24 to get:
----
a = 24(15/16)
---
a = (240+120)/16 = 360/16 = 90/4 = 45/2 = 22.5
----------
Cheers,
Stan H.
-----------